### v1

# On the convergence of noisy Bayesian Optimization with Expected Improvement

Abstract
Expected improvement (EI) is one of the most widely-used acquisition functions in Bayesian optimization (BO). Despite its proven success in applications for decades, important open questions remain on the theoretical convergence behaviors and rates for EI. In this paper, we contribute to the convergence theories of EI in three novel and critical area. First, we consider objective functions that are under the Gaussian process (GP) prior assumption, whereas existing works mostly focus on functions in the reproducing kernel Hilbert space (RKHS). Second, we establish the first asymptotic error bound and its corresponding rate for GP-EI with noisy observations under the GP prior assumption. Third, by investigating the exploration and exploitation of the non-convex EI function, we prove improved error bounds for both the noise-free and noisy cases. The improved noiseless bound is extended to the RKHS assumption as well.
keywords:
Bayesian optimization, expected improvement, convergence rate, Gaussian process prior, noisy observations, non-convex, exploration, exploitation
1 Introduction
Bayesian optimization (BO) is a derivative-free optimization method for black-box functions [1, 2, 3]. BO has seen enormous success in many applications such as machine learning [4], structural design [5], robotics [6], additive manufacturing [7], inertial confinement fusion design [8], etc. In the classic form, BO aims to solve the optimization problem
(1) $\underset{\begin{subarray}{c}{\bm{x}}\end{subarray}\in C}{\text{minimize}}$ $f({\bm{x}}),$
where ${\bm{x}}$ is the decision variable, compact set $C\subset\mathbb{R}^{d}$, and $f:\mathbb{R}^{d}\to\mathbb{R}$ is the objective function.
A surrogate model, often by default a Gaussian process (GP), is used to approximate the black-box objective function (1)  [9]. An acquisition function guides the selection of sequential points for observations. One of the most successful and widely used acquisition functions is the expected improvement (EI) [10]. EI calculates the conditional expectation of an improvement function so that both the mean value and variance of the GP are used in search of the next sample. With a closed form, it is simple to implement in that only the cumulative distribution function (CDF) and probability density function (PDF) of the standard normal distribution are required. Thanks to its effectiveness and efficiency, EI has witnessed rich extensions in literature, e.g., scalable methods [11], constrained EI [12], etc. The classic noise-free BO algorithm using EI is also referred to as the efficient global optimization (EGO) algorithm [13].
Despite its wide adoption and success, existing works on the theoretical convergence of Bayesian optimization with expected improvement, hereby referred to as GP-EI throughout the paper, have clear gaps. The two main streams of research study the asymptotic convergence and the cumulative regret bound of BO algorithms. The asymptotic convergence analysis of GP-EI is limited by the assumptions on the objective and whether observation noise is allowed. Under the assumptions that $f$ is in the reproducing kernel Hilbert space (RKHS) of the GP kernel, also called the frequentist setting, and no observation noise, the asymptotic convergence rates of GP-EI is established in [14], where the convergence rates differ based on the kernels used in the GP. Specifically, define the error $r_{t}^{0}:=f({\bm{x}}_{t}^{+})-f^{*}$, where ${\bm{x}}_{t}^{+}$ is the sample that produces the best current observation among $t$ samples and $f^{*}=\underset{\begin{subarray}{c}{\bm{x}}\in C\end{subarray}}{\text{argmin}}f({\bm{x}})$. The error bound is of $\mathcal{O}(t^{-\frac{1}{d}})$ for squared exponential (SE) kernels and $\mathcal{O}(t^{-\frac{\min\{\nu,1\}}{d}}\log^{\alpha}(t))$ for Matérn kernels, where $\nu$ is the parameter of the kernel and $\alpha=\frac{1}{2}$ for $\nu\in\mathbb{N}$ in Matérn kernels and $\alpha=0$ otherwise. However, to the best of our knowledge, the extensions of [14] to noisy GP-EI or when $f$ follows different assumptions, e.g., $f$ is sampled from a GP, are yet to be accomplished. Other works related to asymptotic convergence of GP-EI include [15], where the authors used no-empty-ball properties of kernels to show that the sequence of iterates generated by GP-EI are dense. In [16], the author showed that for the ranking and selection problems, a modified EI is essentially identical to optimal computing budget allocation (OCBA) algorithm, via the asymptotic sampling ratios. In [17], the author proved consistency of EI for noiseless GP with continuous sample paths.
The other stream of research has focused on the cumulative regret bound, based on the seminal work of [18]. In [19], the authors added additional steps and auxiliary hyper-parameters to modify EI. Further, they assumed bounded conditions in RKHS on the hyper-parameters to derive a cumulative regret bound. In [20], the authors used a stopping criterion to help bound the prediction error. However, the analysis relies on a critical yet not established result that the standard deviation at the optimal point decreases at the same rate as sample points. In [21], the authors attempted to bound the standard deviation at the optimal point via structured initial sampling and achieved an improvement on the regret bound for SE kernels. While effective, their method requires an infinite number of initial samples asymptotically. In [22], the authors proved no-regret for their modified EI, which includes additional control parameters for EI that could become unbounded as $t$ increases.
In most literature mentioned above, $f$ is assumed to be in RKHS of the kernel (see Definition 2.3). Notably however, in [18, 23], the authors analyzed functions $f$ that are sampled from GP priors, also called the Bayesian setting. While the RKHS assumption is critical and general, limitations exist in the functions it represents. Functions in RKHS and functions sampled from GP priors do not encompass each other [18]. In [24], the authors stated that the smoother the kernel, the smaller the RKHS could be for which the convergence theories apply. In cases such as infinite-dimensional Gaussian, [25] calls RKHS small relative to the support of the GP prior it is based on. Additionally, in [23], the authors argued that the constants in RKHS are difficult to compute in practice. Meanwhile, with proper choices of kernels, a function sampled from a GP prior can learn continuous functions with arbitrary precision [26, 25]. In [17], the authors claimed that sample paths of GP often do not satisfy the RKHS assumption. Given its wide adoption in literature, the error bound and convergence analysis of GP-EI where the objective function $f$ is sampled from a GP prior is therefore of utmost importance and interest.
In this paper, we address gaps in the asymptotic convergence analysis of GP-EI mentioned above. Our contributions can be summarized as follows. First, we prove the first asymptotic convergence rate for noisy GP-EI under the GP prior assumption for $f$ (see Assumption 2.4). Given $\delta\in(0,1)$, we prove the upper bound for the noisy error measure $r_{t}:=y^{+}_{t}-f^{*}$, where $y^{+}_{t}$ is the best current observation among $t$ samples, is $\mathcal{O}\left(t^{-\frac{1}{2}}\log(t)^{\frac{d+1}{2}}\right)$ for SE kernels and $\mathcal{O}(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t))$ for Matérn kernels, with probability greater than $1-\delta$. In the noisy GP-EI case, the convergence rate refers to the rate of decrease of the upper bound of $r_{t}$, or equivalently, the rate at which $f({\bm{x}}_{t}^{+})$ converges to the vicinity of the $f^{*}$. Second, we present the asymptotic convergence rates without noise under the GP prior assumption, extending the results in [14] which are obtained under the RKHS assumption. Third, we prove improved error bounds for GP-EI with or without noise under the GP prior assumption. The improved error bounds stem from our novel quantification of the trade-off between exploitation and exploration of EI, the latter being reflected in the posterior standard deviation, at non-sample points. The emerging analysis technique is general and can be used to analyze other convergence measure, e.g., the regret of GP-EI. As an example, we prove a better error bound compared to [14], under the RKHS assumption and noiseless case.
This paper is organized as follows. In Section 2, we describe the basis of the GP-EI algorithm and other background information. In Section 3, we prove multiple properties of EI and the preliminary theoretical results under GP prior assumptions, which are the foundation of the analysis in Section 4. We present our asymptotic convergence analysis on noisy GP-EI under GP prior assumptions in Section 4. In Section 4.4, we improve asymptotic convergence results to noiseless GP-EI under RKHS assumptions. Numerical illustrations of the theories are presented in Section 5, while conclusions are offered in Section 6.
2 Bayesian optimization and expected improvement
There are two main components of Bayesian optimization algorithms: a GP surrogate model for the objective function $f$, and an acquisition function that measures the benefit of the next sample and guides the sequential search. We introduce both in the subsequent subsections.
2.1 Gaussian process
A GP takes the prior distribution of samples to be multivariate Gaussian distribution. Consider a zero mean GP with the kernel (i.e., covariance function) $k({\bm{x}},{\bm{x}}^{\prime}):\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, denoted as $GP(0,k({\bm{x}},{\bm{x}}^{\prime}))$. At each sample ${\bm{x}}_{i}\in C,i=1,2,\dots$, we consider observation noise $\epsilon_{i}$ for $f$ so that the observed function value is $y_{i}=f({\bm{x}}_{i})+\epsilon_{i}$. Thus, the prior distribution on ${\bm{x}}_{1:t}=[{\bm{x}}_{1},\dots,{\bm{x}}_{t}]^{T}$, ${\bm{f}}_{1:t}=[f({\bm{x}}_{1}),\dots,f({\bm{x}}_{t})]^{T}$ and ${\bm{y}}_{1:t}=[y_{1},\dots,y_{t}]^{T}$ is ${\bm{f}}_{1:t}\sim\mathcal{N}(\mathbf{0},{\bm{K}}_{t})$, where $\mathcal{N}$ denotes the normal distribution, and ${\bm{K}}_{t}=[k({\bm{x}}_{1},{\bm{x}}_{1}),\dots,k({\bm{x}}_{1},{\bm{x}}_{t});\dots;k({\bm{x}}_{t},{\bm{x}}_{1}),\dots,k({\bm{x}}_{t},{\bm{x}}_{t})]$.
We assume that the noise follows an independent zero mean Gaussian distribution with variance $\sigma^{2}$, i.e., $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}),i=1,\dots,t$. The posterior distribution of $f({\bm{x}})|{\bm{x}}_{1:t},{\bm{y}}_{1:t}\sim\mathcal{N}(\mu({\bm{x}}),\sigma^{2}({\bm{x}}))$ can be inferred using Bayes’ rule. The posterior mean $\mu_{t}$ and variance $\sigma^{2}_{t}$ are
(2) $\mu_{t}({\bm{x}})\ =\ {\bm{k}}_{t}({\bm{x}})\left({\bm{K}}_{t}+\sigma^{2}{\bm{I}}\right)^{-1}{\bm{y}}_{1:t}$
$\sigma^{2}_{t}({\bm{x}})\ =\ k({\bm{x}},{\bm{x}})-{\bm{k}}_{t}({\bm{x}})^{T}\left({\bm{K}}_{t}+\sigma^{2}{\bm{I}}\right)^{-1}{\bm{k}}_{t}({\bm{x}})\ ,$
where ${\bm{k}}_{t}({\bm{x}})=[k({\bm{x}}_{1},{\bm{x}}),\dots,k({\bm{x}}_{t},{\bm{x}})]^{T}$. When there is no observation noise, (2) can still be used to compute $\mu_{t}$ and $\sigma_{t}$ with a small $\sigma$, e.g., $\sigma=10^{-8}$, for stability of matrix inversion [27, 28].
We make standard assumptions that $k({\bm{x}},{\bm{x}}^{\prime})\leq 1$ and $k({\bm{x}},{\bm{x}})=1$, $\forall{\bm{x}},{\bm{x}}^{\prime}\in C$. Choices of the kernels include the SE function, Matérn functions, etc [9]. The improvement function of $f$ given $t$ samples is defined as
(3) $I_{t}({\bm{x}})=\max\{y^{+}_{t}-f({\bm{x}}),0\},$
where $y^{+}_{t}=\underset{\begin{subarray}{c}y_{i}\in{\bm{y}}_{1:t}\end{subarray}}{\text{argmin}}\ y_{i}$ is the best existing observed objective.
2.2 Expected improvement
In this paper, we consider the probability space $(\Omega,\mathcal{F},\mathbb{P})$, where $\Omega$ is the sample space, $\mathcal{F}$ the $\sigma$-algebra generated by the subspace of $\Omega$ and $\mathbb{P}$ the probability measure on $\mathcal{F}$. We denote a filtration (ordered increasing $\sigma$-algebra) of $\mathcal{F}$ as $\mathcal{F}_{t}$, which is often made equivalent to the history until $t$ defined by: $\mathcal{H}_{t}=\{({\bm{x}}_{s},y_{s}):s=1,\dots,t\}$ in BO literature [29]. The EI acquisition function is defined as the expectation of (3) conditioned on $t$ samples and GP, with a closed form expression:
(4) $EI_{t}({\bm{x}})=(y_{t}^{+}-\mu_{t}({\bm{x}}))\Phi(z_{t}({\bm{x}}))+\sigma_{t}({\bm{x}})\phi(z_{t}({\bm{x}})),$
where
(5) $z_{t}({\bm{x}})=\frac{y^{+}_{t}-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}.$
The functions $\phi$ and $\Phi$ are the PDF and CDF of the standard normal distribution, respectively.
To further analyze the properties of EI, we distinguish between its exploration term, e.g., $\sigma_{t}({\bm{x}})$ in (4), and exploitation term, e.g., $y_{t}^{+}-\mu_{t}({\bm{x}})$ in (4), and define $EI(a,b):\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ as
(6) $EI(a,b)=a\Phi\left(\frac{a}{b}\right)+b\phi\left(\frac{a}{b}\right).$
From the assumptions on kernels in Section 2.1, the exploration $b$ is constrained by $b\in(0,1]$. We view $a$ and $b$ as two independent variables. For a given ${\bm{x}}$, if $a_{t}=y^{+}_{t}-\mu_{t}({\bm{x}})$ and $b_{t}=\sigma_{t}({\bm{x}})\in[0,1]$, we have $EI(a_{t},b_{t})=EI_{t}({\bm{x}})$.
Another commonly used function [21] in the analysis of EI is $\tau:\mathbb{R}\to\mathbb{R}$ defined as
(7) $\tau(z)=z\Phi(z)+\phi(z).$
Thus, EI can be written as $EI_{t}({\bm{x}})=\sigma_{t}({\bm{x}})\tau(z_{t}({\bm{x}}))$. The next sample ${\bm{x}}_{t+1}$ in the GP-EI algorithm is chosen by maximizing the acquisition function over $C$, i.e.,
(8) ${\bm{x}}_{t+1}=\underset{\begin{subarray}{c}x\in C\end{subarray}}{\text{argmax}}EI_{t}({\bm{x}}).$
In order to solve (8), optimization algorithms including L-BFGS or random search can be used. The GP-EI algorithm is given in Algorithm 1, where a stopping criterion can be added.
Algorithm 1 GP-EI algorithm
1:  Choose $\mu_{0}=0$, $k(\cdot,\cdot)$, $\alpha$, and $T_{0}$ initial samples ${\bm{x}}_{i},i=0,\dots,T_{0}$. Observe $y_{i}$.
2:  Train the Gaussian process surrogate model for $f$ on the initial samples.
3:  for $t=T_{0},T_{0}+1,\dots$ do
4:     Find ${\bm{x}}_{t+1}$ based on (8).
5:     Observe $y_{t+1}=f({\bm{x}}_{t+1})+\epsilon_{t+1}$.
6:     Train the surrogate model with the addition of ${\bm{x}}_{t+1}$ and $y_{t+1}$.
7:     if Stopping criterion satisfied then
8:        Exit
2.3 Additional background
We present the remaining necessary background information and our assumptions in this section. In order to find the rate of decrease of the error bound in the noisy case, we use the following well-established lemma for GP based on information theory [18].
Lemma 2.1.
The sum of GP posterior variance $\sigma_{t}({\bm{x}}_{t+1})$ at next sample ${\bm{x}}_{t+1}$ satisfies
(9) $\sum_{i=1}^{t}\sigma_{i-1}^{2}({\bm{x}}_{i})\leq C_{\gamma}\gamma_{t},$
where $C_{\gamma}=\frac{2}{log(1+\sigma^{-2})}$ and $\gamma_{t}$ is the maximum information gain after $t$ samples.
Information gain measures the informativeness of a set of sampling points in $C$ about $f$. Readers are referred to [30, 18] for a detailed definition of the maximum information gain $\gamma_{t}>0$. The rate of increase for $\gamma_{t}$ is dependent on the property of the kernel. For common kernels such as the SE kernel and the Matérn kernel, $\gamma_{t}$ and its order of increase have been widely studied in literature [31]. The state-of-the-art rates of $\gamma_{t}$ for two commonly-used kernels are summarized below [31].
Lemma 2.2.
For Gaussian process with $t$ samples, the SE kernel has
$\gamma_{t}=\mathcal{O}(\log^{d+1}(t))$, and the Matérn kernel with smoothness parameter $\nu>0$ has $\gamma_{t}=\mathcal{O}(t^{\frac{d}{2\nu+d}}(\log^{\frac{2\nu}{2\nu+d}}(t)))$.
The formal definition of RKHS is given below.
Definition 2.3.
Consider a positive definite kernel $k:C\times C\to\mathbb{R}$ with respect to a finite Borel measure supported on $C$. A Hilbert space $H_{k}$ of functions on $C$ with an inner product $\langle\cdot,\cdot\rangle_{H_{k}}$ is called a RKHS with kernel $k$ if for $\forall{\bm{x}}\in C$, $k(\cdot,{\bm{x}})\in H_{k}$, and $\forall{\bm{x}}\in C,f\in H_{k}$, $\langle f,k(\cdot,{\bm{x}})\rangle_{H_{k}}=f({\bm{x}})$. The induced RKHS norm $\left\lVert{f}\right\rVert_{H_{k}}=\sqrt{\langle f,f\rangle_{H_{k}}}$ measures the smoothness of $f$ with respect to $k$.
For ease of reference, we list the assumptions used in our analysis below, all of which are common in BO literature. First, we have the GP prior assumption for $f$, also called the Bayesian setting in literature.
Assumption 2.4.
The objective function $f$ is a sample from the Gaussian process $GP(0,k({\bm{x}},{\bm{x}}^{\prime}))$, where $k({\bm{x}},{\bm{x}}^{\prime})\leq 1$ and $k({\bm{x}},{\bm{x}})=1$. Further, $f$ is Lipschitz continuous with constant $L$.
The RKHS assumption of $f$ is given below, also called the frequentist setting.
Assumption 2.5.
The objective function $f$ is in the RKHS $\mathcal{H}_{k}$ associated with the kernel $k({\bm{x}},{\bm{x}}^{\prime})\leq 1$ and $k({\bm{x}},{\bm{x}})=1$. Further, the RKHS norm of $f$ is bounded $\left\lVert{f}\right\rVert_{\mathcal{H}_{k}}\leq B$ for some $B\geq 1$.
Without losing generality, the set constraint assumption is given below.
Assumption 2.6.
The set $C\subseteq[0,r]^{d}$ is compact for $r>0$.
The Gaussian assumption on the noise is given below.
Assumption 2.7.
The observation noise are i.i.d. random samples from a zero mean Gaussian, i.e., $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}),\sigma>0$ for all $t$.
3 Preliminary results
We present the properties of the non-convex EI function in Section 3.1 and the preliminary lemmas for GP-EI under the GP prior assumption in Section 3.2, which are both important for the convergence analysis in Section 4. We use union bound repeatedly to obtain the probability of multiple events or inequalities occurring. Without losing generality, we consider the case $\sigma_{t}({\bm{x}})>0$, as the results for the case $\sigma_{t}({\bm{x}})=0$ hold trivially.
3.1 EI properties
First, we state some basic properties of $\phi$, $\Phi$, and $\tau$ in the following lemma.
Lemma 3.1.
The PDF and CDF of standard normal distribution satisfy $0<\phi(x)\leq\phi(0),\Phi(x)\in(0,1)$, for any $x\in\mathbb{R}$. Given a random variable $\xi$ sampled from the standard normal distribution, i.e., $\xi\sim\mathcal{N}(0,1)$, we have $\mathbb{P}\{\xi>c|c>0\}\leq\frac{1}{2}e^{-c^{2}/2}$. Similarly, for $c<0$, $\mathbb{P}\{\xi<c|c<0\}\leq\frac{1}{2}e^{-c^{2}/2}$.
The last statement in Lemma 3.1 is a well-known result (e.g., see proof of Lemma 5.1 in [18]), as Figure 1 (left) demonstrates. The property of $\tau(\cdot)$ in (7) is given below.
Lemma 3.2.
The function $\tau(\cdot)$ is monotonically increasing and $\tau(z)>0$ for $\forall z\in\mathbb{R}$. Moreover, $\frac{d\tau(z)}{dz}=\Phi(z)$.
Proof 3.3.
From the definition of $\tau(z)$, we can write
[…]
Given the definition of $\phi(u)$,
(11) $\frac{d\phi(u)}{du}=\frac{1}{\sqrt{2\pi}}e^{-\frac{u^{2}}{2}}(-u)=-\phi(u)u.$
Thus, the derivative of $\tau$ is
(12) $\frac{d\tau(z)}{dz}=\Phi(z)+z\phi(z)-\phi(z)z=\Phi(z)>0,$
which completes the proof.
Another lemma for $\tau(\cdot)$ is given below, with an illustrative plot in Figure 1 (right).
Lemma 3.4.
Given $z>0$, $\Phi(-z)>\tau(-z)$.
Proof 3.5.
Define $q(z)=\Phi(-z)-\tau(-z)$. Using integration by parts, we have
[…]
Replacing $z$ with $-z$ in (13),
(14) $\phi(-z)\left(\frac{1}{z}-\frac{1}{z^{3}}\right)<\Phi(-z).$
Multiplying both sides in (14) by $1+z$,
(15) $(1+z)\Phi(-z)>\phi(-z)\frac{z^{2}-1}{z^{3}}(1+z)=\phi(-z)\left(1+\frac{z^{2}-z-1}{z^{3}}\right).$
Thus, if $z>\frac{1+\sqrt{5}}{2}$, then the right-hand-side of (15) is greater than $\phi(-z)$ and
(16) $q(z):=\Phi(-z)-\tau(-z)=(1+z)\Phi(-z)-\phi(-z)>0.$
Consequently, we focus on $z\in(0,\frac{1+\sqrt{5}}{2}]$. Taking the derivative of $q(z)$, by Lemma 3.2,
(17) $q^{\prime}(z):=\frac{dq(z)}{dz}=-\phi(-z)+\Phi(-z).$
Further, the derivative of $q^{\prime}(z)$ is
(18) $\frac{d^{2}q(z)}{dz^{2}}=\frac{dq^{\prime}(z)}{dz}=-\phi(-z)+\phi(-z)z=\phi(z)(z-1).$
For $z>1$, $\frac{d^{2}q(z)}{dz^{2}}>0$. For $z\in(0,1)$ $\frac{d^{2}q(z)}{dz^{2}}<0$. Thus, $q^{\prime}(z)$ is monotonically increasing for $z>1$ and decreasing for $z\in(0,1)$. For $z>1$, from Lemma 3.2,
(19) $q^{\prime}(z)=\Phi(-z)-\phi(-z)<z\Phi(-z)-\phi(-z)=-\tau(-z)<0.$
Further, $q^{\prime}(0)=\Phi(0)-\phi(0)>0$ and $q^{\prime}(1)=\Phi(-1)-\phi(-1)<0$. Thus, there exists unique $\bar{z}\in(0,1)$ so that $q^{\prime}(\bar{z})=0$. For $z\in(0,\bar{z})$, $q(z)$ is monotonically increasing. For $z\in[\bar{z},\frac{1+\sqrt{5}}{2}]$, $q(z)$ is monotonically decreasing. Therefore, $q(z)>\min\{q(0),q(\frac{1+\sqrt{5}}{2})\}$ for $z\in(0,\frac{1+\sqrt{5}}{2})$. Since $q(0)>0$ and $q(\frac{1+\sqrt{5}}{2})>0$, $q(z)>0$ for $z\in(0,\frac{1+\sqrt{5}}{2})$. Combined with (16), the proof is complete.
The next lemma proves basic properties for $EI_{t}$.
Lemma 3.6.
For $\forall{\bm{x}}\in C$, $EI_{t}({\bm{x}})\geq 0$ and $EI_{t}({\bm{x}})\geq y^{+}_{t}-\mu_{t}({\bm{x}})$. Moreover,
(20) $z_{t}({\bm{x}})\leq\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}<\begin{cases}\phi(z_{t}({\bm{x}})),\ &z_{t}({\bm{x}})<0\\ z_{t}({\bm{x}})+\phi(z_{t}({\bm{x}})),\ &z_{t}({\bm{x}})\geq 0.\end{cases}$
Proof 3.7.
From the definition of $I_{t}$ and $EI_{t}$, the first statement follows immediately. By (4),
(21) $\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}=z_{t}({\bm{x}})\Phi(z_{t}({\bm{x}}))+\phi(z_{t}({\bm{x}})).$
If $z_{t}({\bm{x}})<0$, or equivalently $y^{+}_{t}-\mu_{t}({\bm{x}})<0$, (21) leads to $\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}<\phi(z_{t}({\bm{x}}))$. If $z_{t}({\bm{x}})\geq 0$, $\Phi(\cdot)<1$ gives us $\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}<z_{t}({\bm{x}})+\phi(z_{t}({\bm{x}}))$. The left inequality in (20) is a direct result of $EI_{t}({\bm{x}})\geq y^{+}_{t}-\mu_{t}({\bm{x}})$.
The monotonicity of the exploration and exploitation of (6) is given next.
Lemma 3.8.
$EI(a,b)$ is monotonically increasing with respect to $a$ and $b$ for $b\in(0,1]$.
Proof 3.9.
Taking the derivative of $EI(a,b)$ with respect to $a$,
(22) $\frac{\partial EI(a,b)}{\partial a}=\Phi\left(\frac{a}{b}\right)+a\phi\left(\frac{a}{b}\right)\frac{1}{b}+b\frac{\partial\phi\left(\frac{a}{b}\right)}{\partial a}.$
From (11), (22) is
(23) $\frac{\partial EI(a,b)}{\partial a}=\Phi\left(\frac{a}{b}\right)+\phi\left(\frac{a}{b}\right)\frac{a}{b}-\phi\left(\frac{a}{b}\right)\frac{a}{b}=\Phi\left(\frac{a}{b}\right)>0.$
Similarly,
(24) $\frac{\partial EI(a,b)}{\partial b}=$ $-a\phi\left(\frac{a}{b}\right)\frac{a}{b^{2}}+\phi\left(\frac{a}{b}\right)-b\phi\left(\frac{a}{b}\right)\frac{a}{b}(-\frac{a}{b^{2}})=\phi\left(\frac{a}{b}\right)>0.$
3.2 Characterizations of GP prior objective
The following lemma is on the cumulative probability distribution of $I_{t}(\cdot)$ under the GP prior assumption.
Lemma 3.10.
Under Assumption 2.4, the cumulative probability distribution of $I_{t}$ satisfies
(25) $\mathbb{P}\{I_{t}({\bm{x}})\leq a\}=\begin{cases}0,\ &a<0,\\ \Phi\left(\frac{a}{\sigma_{t}({\bm{x}})}-z_{t}({\bm{x}})\right),\ &a\geq 0.\end{cases}$
Proof 3.11.
Under Assumption 2.4, at a given $t$, $f({\bm{x}})\sim\mathcal{N}(\mu_{t}({\bm{x}}),\sigma_{t}({\bm{x}}))$. Since $I_{t}({\bm{x}})\geq 0$ for all ${\bm{x}}$, (25) follows immediately if $a<0$. For $a\geq 0$,
(26) $\mathbb{P}\{I_{t}({\bm{x}})\leq a\}=\mathbb{P}\{y^{+}_{t}-f({\bm{x}})\leq a\}=1-\mathbb{P}\{f({\bm{x}})\leq y^{+}_{t}-a\}.$
Using basic properties of the CDF of the standard normal distribution,
(27) $1-\mathbb{P}\{f({\bm{x}})\leq y^{+}_{t}-a\}=1-\Phi\left(\frac{y_{t}^{+}-a-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right)=\Phi\left(\frac{a-y_{t}^{+}+\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right).$
The next lemma is a well-known result on the prediction error $f({\bm{x}})-\mu_{t}({\bm{x}})$ under Assumption 2.4 [18].
Lemma 3.12.
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{1}{\delta})$. Under Assumption 2.4, for any given ${\bm{x}}\in C$ and $t\in\mathbb{N}$, the following holds
(28) $\mathbb{P}\{|f({\bm{x}})-\mu_{t}({\bm{x}})|\leq\sqrt{\beta}\sigma_{t}({\bm{x}})\}\geq 1-\delta.$
Moreover, the one-sided inequalities hold with probability greater than $1-\frac{\delta}{2}$, i.e., $\mathbb{P}\{f({\bm{x}})-\mu_{t}({\bm{x}})\leq\sqrt{\beta}\sigma_{t}({\bm{x}})\}\geq 1-\frac{\delta}{2}$, and $\mathbb{P}\{f({\bm{x}})-\mu_{t}({\bm{x}})\geq-\sqrt{\beta}\sigma_{t}({\bm{x}})\}\geq 1-\frac{\delta}{2}$.
Proof 3.13.
Under Assumption 2.4, $f({\bm{x}})\sim\mathcal{N}(\mu_{t}({\bm{x}}),\sigma_{t}^{2}({\bm{x}}))$. By Lemma 3.1,
(29) $\mathbb{P}\left\{f({\bm{x}})-\mu_{t}({\bm{x}})>\sqrt{\beta}\sigma_{t}({\bm{x}})\right\}=1-\Phi\left(\sqrt{\beta}\right)\leq\frac{1}{2}e^{-\frac{\beta}{2}}.$
Similarly,
(30) $\mathbb{P}\left\{f({\bm{x}})-\mu_{t}({\bm{x}})<-\sqrt{\beta}\sigma_{t}({\bm{x}})\right\}\leq\frac{1}{2}e^{-\frac{\beta}{2}}.$
Thus,
(31) $\mathbb{P}\left\{|f({\bm{x}})-\mu_{t}({\bm{x}})|<\sqrt{\beta}\sigma_{t}({\bm{x}})\right\}\geq 1-e^{-\frac{\beta}{2}}.$
Let $\beta$ be such that $e^{-\frac{\beta}{2}}=\delta$ and (28) is proven. Similarly, from (29) and (30), we obtain the one-sided inequalities.
Lemma 3.12 is extended to $\forall t$ and ${\bm{x}}_{t}$ using union bound and an increasing $\beta_{t}$ in the following lemma (see [18] for proof).
Lemma 3.14.
Given $\delta\in(0,1)$, let $\beta_{t}=2\log(\frac{\pi_{t}}{\delta})$, where $\pi_{t}=\frac{\pi^{2}t^{2}}{6}$. Under Assumption 2.4, for $\forall t\in\mathbb{N}$, the following holds
(32) $\mathbb{P}\{|f({\bm{x}}_{t+1})-\mu_{t}({\bm{x}}_{t+1})|\leq\sqrt{\beta_{t+1}}\sigma_{t}({\bm{x}}_{t+1}),\forall t\in\mathbb{N}\}\geq 1-\delta.$
Next, we establish the relationship between $I_{t}$ and $EI_{t}({\bm{x}})$.
Lemma 3.15.
Given $\delta\in(0,1)$, let $\beta=\max\{1.44,2\log(\frac{c_{\alpha}}{\delta})\}$, where constant $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Under Assumption 2.4, at given ${\bm{x}}\in C$ and $t\geq\mathbb{N}$,
(33) $\mathbb{P}\left\{|I_{t}({\bm{x}})-EI_{t}({\bm{x}})|\leq\sqrt{\beta}\sigma_{t}({\bm{x}})\right\}\geq 1-\delta.$
Proof 3.16.
Given a scalar $w>1$, we consider the probabilities
(34) $\mathbb{P}\left\{I_{t}({\bm{x}})>\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\right\}\quad\text{and}\quad\mathbb{P}\left\{I_{t}({\bm{x}})<-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\right\}.$
Consider the first probability of (34). From Lemma 3.6, $EI_{t}({\bm{x}})\geq 0$ for $\forall{\bm{x}}$ and $t$. Therefore, $\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})>0$. From Lemma 3.6, Lemma 3.10, and the monotonicity of $\Phi$, we have
(35) $\mathbb{P}\left\{I_{t}({\bm{x}})>\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\right\}=$ $1-\Phi\left(\frac{\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})-y^{+}_{t}+\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right)$
$\leq$ $1-\Phi(w)\leq\frac{1}{2}e^{-\frac{w^{2}}{2}},$
where the last inequality is from Lemma 3.1.
For the second probability in (34), we further distinguish between two cases. First, consider $-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})<0$. From Lemma 3.10,
(36) $\mathbb{P}\left\{I_{t}({\bm{x}})<-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\right\}=0.$
Second, $-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\geq 0$. By Lemma 3.10, we have
(37) $\mathbb{P}\left\{I_{t}({\bm{x}})<-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\right\}=\Phi\left(-w+\frac{EI_{t}({\bm{x}})-y^{+}_{t}+\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right).$
To proceed, we show that $y^{+}_{t}-\mu_{t}({\bm{x}})\geq 0$. Suppose on the contrary, $y^{+}_{t}-\mu_{t}({\bm{x}})<0$ and thus $z_{t}({\bm{x}})<0$. From Lemma 3.6,
(38) $\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}<\phi(z_{t}({\bm{x}}))\leq\phi(0)<1\leq w.$
Then, (38) contradicts the premise of this case. Thus, we have $y^{+}_{t}-\mu_{t}({\bm{x}})\geq 0$ and $z_{t}({\bm{x}})\geq 0$. From the definition (4) and $\Phi\in(0,1)$,
(39) $\frac{EI_{t}({\bm{x}})-y^{+}_{t}+\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}=$ $\left[z_{t}({\bm{x}})\left(\Phi(z_{t}({\bm{x}}))-1\right)+\phi(z_{t}({\bm{x}}))\right]<\phi(z_{t}({\bm{x}})).$
In addition, by the premise of this case and Lemma 3.6,
(40) $w\leq\frac{EI_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\leq z_{t}({\bm{x}})+\phi(z_{t}({\bm{x}})).$
Given that $w>1$ and $\phi(0)\geq\phi(z_{t}({\bm{x}}))$, we have
(41) $z_{t}({\bm{x}})+\phi(0)>z_{t}({\bm{x}})+\phi(z_{t}({\bm{x}}))>w,\ z_{t}({\bm{x}})>w-\phi(0)>0.$
As $z_{t}({\bm{x}})\geq 0$ increases, $\phi(z_{t}({\bm{x}}))>0$ decreases. Thus, we have
(42) $\frac{z_{t}({\bm{x}})}{\phi(z_{t}({\bm{x}}))}>\frac{w-\phi(0)}{\phi(w-\phi(0))},\ \phi(z_{t}({\bm{x}}))<\frac{\phi(w-\phi(0))}{w-\phi(0)}z_{t}({\bm{x}}).$
Denote $c_{1}(w)=\frac{w-\phi(0)}{w-\phi(0)+\phi(w-\phi(0))}$. Applying (42) to (40), we obtain
(43) $c_{1}(w)w<z_{t}({\bm{x}}),\ \ \phi(z_{t}({\bm{x}}))<\phi(c_{1}(w)w).$
Applying (43) and (39) to (37), we obtain
(44) $\mathbb{P}\left\{I_{t}({\bm{x}})<-w\sigma_{t}({\bm{x}})+EI_{t}({\bm{x}})\right\}<$ $\Phi\left(-w+\phi(z_{t}({\bm{x}}))\right)$
$<$ $\Phi\left(-w+\phi(c_{1}(w)w)\right).$
Notice that $\phi(c_{1}(w)w)<\phi(c_{1}(w))<\phi(c_{1}(w))w$ due to $w>1$. By the definition of $\Phi$ and mean value theorem,
[…]
where $c_{2}(w)=[1-\phi(c_{1}(w))]^{2}+[c_{1}(w)]^{2}$. The last inequality in (45) again uses Lemma 3.1. Notice that $c_{2}(w)$ increases with $w$ and for $w\geq 1.2$, $c_{2}(w)>11$. Thus, $e^{-\frac{1}{2}w^{2}}>e^{-\frac{1}{2}c_{2}(w)w^{2}}$ for $w\geq 1.2$, which simplifies (45) to
(46) $\Phi\bigl{(}-w+$ $\phi(c_{1}(w)w)\bigr{)}$ $<\frac{1+\pi}{2\pi}e^{-\frac{1}{2}w^{2}}.$
Therefore, by (44) and (46), if $w\geq 1.2$,
(47) $\mathbb{P}\left\{I_{t}({\bm{x}})<-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\right\}<\frac{1+\pi}{2\pi}e^{-\frac{1}{2}w^{2}}.$
Combining (47) with (35) and (36), we have
(48) $\mathbb{P}\left\{\left|I_{t}({\bm{x}})-EI_{t}({\bm{x}})\right|>w\sigma_{t}({\bm{x}})\right\}<\frac{1+2\pi}{2\pi}e^{-\frac{1}{2}w^{2}},$
for $w\geq 1.2$. The probability in (48) monotonically decreases with $w$. Let
(49) $\delta=\frac{1+2\pi}{2\pi}e^{-\frac{1}{2}w^{2}}.$
Taking the logarithm of (49) leads to $\log(\frac{1+2\pi}{2\pi\delta})=\frac{1}{2}w^{2}$. Let $\beta=\max\{w^{2},1.2^{2}\}$, and the proof is complete.
Another relationship between $I_{t}$ and $EI_{t}$ under the GP prior assumption is given in the following lemma.
Lemma 3.17.
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{1}{\delta})$. Under Assumption 2.4, at given ${\bm{x}}\in C$ and $t\geq 1$,
(50) $\mathbb{P}\left\{\frac{\tau(-\sqrt{\beta})}{\tau(\sqrt{\beta})}I_{t}({\bm{x}})\leq EI_{t}({\bm{x}})\right\}\geq 1-\delta.$
Proof 3.18.
From Lemma 3.12, with probability $\geq 1-\delta$, (28) stands. If $y^{+}_{t}-f({\bm{x}})\leq 0$, then $I_{t}({\bm{x}})=0$. Since $EI_{t}({\bm{x}})\geq 0$, (50) is trivial. If $y^{+}_{t}-f({\bm{x}})>0$, by Lemma 3.12,
(51) $y^{+}_{t}-\mu_{t}({\bm{x}})$ $=y^{+}_{t}-f({\bm{x}})+f({\bm{x}})-\mu_{t}({\bm{x}})>f({\bm{x}})-\mu_{t}({\bm{x}})$
$>-\sqrt{\beta}\sigma_{t}({\bm{x}}),$
with probability greater than $\geq 1-\frac{\delta}{2}$. From Lemma 3.2, we have
(52) $\tau\left(\frac{y^{+}_{t}-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right)>\tau(-\sqrt{\beta}),$
and therefore,
(53) $EI_{t}({\bm{x}})=\sigma_{t}({\bm{x}})\tau\left(\frac{y^{+}_{t}-\mu_{t}({\bm{x}})}{\sigma_{t}({\bm{x}})}\right)>\tau(-\sqrt{\beta})\sigma_{t}({\bm{x}}),$
with probability greater than $1-\frac{\delta}{2}$. Next, we let $w=\sqrt{\beta}$ and consider the first probability in (34). Since $\sqrt{\beta}\sigma_{t}({\bm{x}})+EI_{t}({\bm{x}})>0$, we can follow the proof procedure and reach (35) so that
(54) $\mathbb{P}\left\{I_{t}({\bm{x}})>\sqrt{\beta}\sigma_{t}({\bm{x}})+EI_{t}({\bm{x}})\right\}\leq\frac{1}{2}e^{-\frac{\beta}{2}}=\frac{\delta}{2}.$
Therefore,
(55) $\mathbb{P}\left\{I_{t}({\bm{x}})-EI_{t}({\bm{x}})\leq\sqrt{\beta}\sigma_{t}({\bm{x}})\right\}\geq 1-\frac{\delta}{2}.$
Applying (55) to (53) with union bound leads to
(56) $EI_{t}({\bm{x}})>\frac{\tau(-\sqrt{\beta})}{\sqrt{\beta}+\tau(-\sqrt{\beta})}I_{t}({\bm{x}})=\frac{\tau(-\sqrt{\beta})}{\tau(\sqrt{\beta})}I_{t}({\bm{x}}),$
with probability greater than $1-\delta$.
4 Convergence analysis for GP-EI
In this section, we consider the asymptotic convergence of GP-EI under Assumption 2.4. Since $y^{+}_{t}$ is the actual output of GP-EI and plays an integral part in the algorithm itself (Algorithm 1), we consider the error measure $r_{t}$ given $t$ samples defined by
(57) $r_{t}=y^{+}_{t}-f({\bm{x}}^{*}),$
where $f({\bm{x}}^{*})=f^{*}$. The error (57) is a natural extension of the noiseless error $f({\bm{x}}^{+}_{t})-f({\bm{x}}^{*})$ in [14], which is also called simple regret. Hence, the asymptotic convergence in the noisy case refers to the best existing observation converging (with probability $1-\delta$ for $\forall\delta\in(0,1)$) to the vicinity of $f^{*}$ with values no greater than $f^{*}$, i.e., $r_{t}\leq 0$ as $t\to\infty$.
While (57) is not guaranteed to be positive given the nature of noisy observations, $r_{t}$ is monotonically non-increasing, i.e., $r_{t+1}\leq r_{t}$. More importantly, our asymptotic convergence analysis of $r_{t}$ can be easily used to establish convergence of various error measures under additional conditions. For instance, a lower bound of (57) in terms of $\sigma_{t}({\bm{x}}_{t+1})$ can be established if a stopping criterion such as $EI_{t}({\bm{x}}_{t+1})\geq\kappa,\kappa>0$ is used, similar to [20]. Then, $r_{t}\to 0$ can be achieved if the stopping criterion is not met for $\forall t\in\mathbb{N}$. Equivalently, one can adopt a different error measure
(58) $\tilde{r}_{t}=f({\bm{x}}_{t})-f({\bm{x}}^{*})\geq 0,$
and obtain convergence with the stopping criterion. Another error measure for GP-EI is
(59) $r_{t}^{\prime}=I_{t}({\bm{x}}^{*})=\max\{y^{+}_{t}-f({\bm{x}}^{*}),0\}\geq 0,$
which measures the error at $t$ against ${\bm{x}}^{*}$ in terms of the improvement function $I_{t}$, an integral part of the GP-EI algorithm. It is clear that $r_{t}\leq 0$ as $t\to\infty$ directly leads to the convergence of (59).
Assumption 2.42.6 and 2.7 are used in this section. In most of our analysis, we present error bound in terms of $\sigma_{t}({\bm{x}}_{t+1})$, which satisfies $\sigma_{t}({\bm{x}}_{t+1})\to 0$ as $t\to\infty$, for many commonly-used kernels and GP-based algorithms [18, 14]. In the noiseless case, the salient idea is that $\sigma_{t}({\bm{x}}_{t+1})$ can be bounded by the maximum distance between existing $t$ samples and any point $\forall{\bm{x}}\in C$, leading to convergence rate results for specific kernels. To improve noiseless error bound, we adopt the kernel assumptions in [14] and provide the convergence rates in Theorem 4.9. However, the techniques do not apply to the noisy case. Instead, we prove convergence rates in the noisy case using information theory [18] for two of the most widely used kernels: the SE kernel and Matérn kernel in Theorem 4.12.
For clarity of presentation, we briefly summarize the results for this section. Theorem 4.2 establishes the asymptotic error bounds with and without noise under GP prior assumptions. We point out that the noiseless error bound is similar to that in [14] under the RKHS assumption with no noise. Theorem 4.7 takes advantage of the exploration and exploitation properties of EI to establish a better error bound than Theorem 4.2 with and without noise (under the GP prior assumption). In Theorem 4.12, the convergence rates in the noisy case using the improved error bounds from Theorem 4.7 are given for SE and Matérn kernels. In Proposition 4.14, we demonstrate the improvement of Theorem 4.7 compared to Theorem 4.2.
4.1 Asymptotic convergence in the noisy case
We state the boundedness of $f$ on $C$ as a lemma for easy reference.
Lemma 4.1.
There exists $M>0$, such that $|f({\bm{x}})|\leq M$ for all ${\bm{x}}\in C$.
Theorem 4.2.
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{6}{\delta})$ and $c_{t}^{\sigma}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$ The error bound of the EI algorithm satisfies
(60) $\mathbb{P}\left\{r_{t}\leq c_{\tau}(\beta)\left[3\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{t-3}+(\sqrt{\beta}+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right]\right\}\geq 1-\delta,$
for given $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, where $c_{\tau}(\beta)=\frac{\tau(\sqrt{\beta})}{\tau(-\sqrt{\beta})}$ and $t_{k}\in[\frac{t}{3}-1,t]$. If there is no observation noise, i.e., $\sigma=0$, let $\beta=2\log(\frac{2}{\delta})$. Then,
(61) $\mathbb{P}\left\{r_{t}\leq c_{\tau}(\beta)\left[2\frac{2M}{t-2}+(\sqrt{\beta}+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right]\right\}\geq 1-\delta,$
for $t_{k}\in[\frac{t}{2}-1,t]$.
Proof 4.3.
Since $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$, by Lemma 3.1, $\mathbb{P}\left\{\epsilon_{t}\leq\sqrt{w}\sigma\right\}=1-\Phi(-\sqrt{w})\geq 1-\frac{1}{2}e^{-\frac{1}{2}w}$, for any given $t\in\mathbb{N}$, where $w>0$ is a scalar. Similar to Lemma 3.14, if $w_{t}=2\log(\frac{\pi^{2}t^{2}}{6\delta})$, we can use union bound and write $\mathbb{P}\left\{|\epsilon_{t}|\leq\sqrt{w_{t}}\sigma\right\}\geq 1-\delta$, for all $t\in\mathbb{N}$. Let $c^{\sigma}_{t}=w_{t}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$ so that
(62) $\mathbb{P}\left\{|\epsilon_{t}|\leq\sqrt{c_{t}^{\sigma}}\sigma\right\}\geq 1-\frac{\delta}{3},$
for all $t\in\mathbb{N}$. From Lemma 4.1, (62), and union bound, with probability greater than $1-\frac{\delta}{3}$,
(63) $\sum_{t=0}^{T-1}y^{+}_{t}-y^{+}_{t+1}=y^{+}_{0}-y_{T}^{+}=f({\bm{x}})+\epsilon_{0}-f({\bm{x}}^{+}_{T})-\epsilon_{T}^{+}\leq 2M+2\sqrt{c^{\sigma}_{t}}\sigma.$
Given that $y^{+}_{t}-y^{+}_{t+1}\geq 0$, $y^{+}_{t}-y^{+}_{t+1}\geq\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k}$ at most $k$ times for any $k\in\mathbb{N}$ with probability greater than $1-\frac{\delta}{3}$.
Consider the observation noise with $\sigma>0$. Let $A_{k}$ be the index set for the first $k$ samples, where $|A_{k}|=k$, $k\in\mathbb{N}$. For a sequence of iterates ${\bm{x}}_{i},i\in A_{k}$, the probability of $f({\bm{x}}_{i})<y_{i}=f({\bm{x}}_{i})+\epsilon_{i}$ for all $i\in A_{k}$ is $(\frac{1}{2})^{k}$, thanks to the i.i.d. $\epsilon_{i}$. Equivalently, the probability that there exists $i\in A_{k}$ such that $f({\bm{x}}_{i})\geq y_{i}$ is $1-(\frac{1}{2})^{k}$. Let $k=[\frac{t}{3}]$ where $[x]$ is the largest integer smaller than $x$. Then, $3k\leq t\leq 3(k+1)$. For $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, $(\frac{1}{2})^{k}\leq(\frac{1}{2})^{\frac{t}{3}-1}\leq\frac{\delta}{3}$. Thus, using union bound, there exists $k\leq t_{k}\leq 3k$ so that
(64) $y_{t_{k}}^{+}-y_{t_{k}+1}^{+}<\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k},\ f({\bm{x}}_{t_{k}+1})\geq y_{t_{k}+1}\geq y_{t_{k}+1}^{+},$
with probability greater than $1-\frac{2\delta}{3}$ for $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$.
The choice of $\beta$ satisfies both Lemma 3.12 and 3.17 with $\frac{\delta}{6}$ each. Using $\frac{\delta}{6}$ and $t_{k}$ in Lemma 3.17, with probability greater than $1-\frac{\delta}{6}$,
(65) $r_{t}=$ $y^{+}_{t}-f({\bm{x}}^{*})\leq y^{+}_{t_{k}}-f({\bm{x}}^{*})=I_{t_{k}}({\bm{x}}^{*})$
$\leq$ $\frac{\tau(\sqrt{\beta})}{\tau(-\sqrt{\beta})}EI_{t_{k}}({\bm{x}}^{*})\leq\frac{\tau(\sqrt{\beta})}{\tau(-\sqrt{\beta})}EI_{t_{k}}({\bm{x}}_{t_{k}+1})$
$=$ $c_{\tau}(\beta)\left[(y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1}))\Phi(z_{t_{k}}({\bm{x}}_{t_{k}+1}))+\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\phi(z_{t_{k}}({\bm{x}}_{t_{k}+1}))\right].$
Using $\frac{\delta}{6}$ and $t_{k}$ in Lemma 3.12, with probability greater than $1-\frac{5\delta}{6}$,
(66) $y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})=$ $y^{+}_{t_{k}}-y^{+}_{t_{k}+1}+y_{t_{k}+1}^{+}-f({\bm{x}}_{t_{k}+1})+f({\bm{x}}_{t_{k}+1})-\mu_{t_{k}}({\bm{x}}_{t_{k}+1}),$
$\leq$ $\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k}+\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}),$
for $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$. Applying (66) to (65) leads to
(67) $\mathbb{P}\left\{r_{t}\leq c_{\tau}({\beta})\left[\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k}+(\sqrt{\beta}+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right]\right\}\geq 1-\delta,$
for $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, where $\phi(\cdot)\leq\phi(0)$ is used.
Next, consider the noiseless case. Choose $k=[\frac{t}{2}]$ so that $2k\leq t\leq 2(k+1)$. There exists $k\leq t_{k}\leq 2k$ so that $y^{+}_{t_{k}}-y^{+}_{t_{k}+1}<\frac{2M}{k}$. Using $\frac{\delta}{2}$ and $t_{k}$ in Lemma 3.17, we have with probability greater than $1-\frac{\delta}{2}$, (65) stands. Given no noise, $y^{+}_{t}\leq f({\bm{x}}_{t})$ for all $t\in\mathbb{N}$. Using $\frac{\delta}{2}$ and $t_{k}$ in Lemma 3.12, we have
(68) $\centering y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})=$ $y^{+}_{t_{k}}-y^{+}_{t_{k}+1}+y^{+}_{t_{k}+1}-f(x_{t_{k}+1})+f(x_{t_{k}+1})-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})$
$<$ $\frac{2M}{k}+f({\bm{x}}_{t_{k}+1})-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})\leq\frac{2M}{k}+\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}),$
with probability greater than $1-\frac{\delta}{2}$. From (68) and Lemma 3.1, (65) simplifies to
(69) $\mathbb{P}\left\{r_{t}\leq c_{\tau}({\beta})\left[\frac{2M}{k}+(\sqrt{\beta}+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right]\right\}\geq 1-\delta.$
Remark 4.4.
The constants $c_{\tau}(\beta)$ is dependent on $\delta$ and increases quickly as $\delta$ decreases. The step number $t_{k}$ increases with $t$. Therefore, if $\sigma_{t}({\bm{x}}_{t+1})\to 0$, $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\to 0$.
Remark 4.5.
The noiseless case in Theorem 4.2 is similar to Theorem 2 in [14], which assumes the objective in RKHS. Using the kernel assumptions (1)-(4) on $k({\bm{x}},{\bm{x}}^{\prime})$ in [14], Theorem 4.2 leads to the $k^{-\frac{\min\{\nu,1\}}{d}}\log^{\beta}(k)$ rate for $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$ where $k\leq t_{k}\leq 3k$, and the same rate for $r_{t}$ as in [14]. We formalize the rates in the noiseless case in Theorem 4.9. However, these rates do not hold in the noisy case as $\sigma_{t}$ does not follow Lemma 7 in [14] anymore. The noisy case, where additional uncertainty is introduced into GP-EI, generates a larger upper bound with a larger $t$ for the same probability. We prove convergence rates for two kernels using information theory in the noisy case in Theorem 4.12.
Remark 4.6.
From a technical perspective, two main challenges existed for the asymptotic convergence extension of [14]. The first one is the lack of a pointwise upper bound in the order of $\sigma_{t}({\bm{x}})$ on the prediction error $f({\bm{x}})-\mu_{t}({\bm{x}})$ when $f$ is in RKHS. For instance, if $f$ is in RKHS with sub-Gaussian noises $\epsilon_{i},i=1,\dots,t$[29] proves $|f({\bm{x}})-\mu_{t}({\bm{x}})|\leq\sqrt{\beta_{t}}\sigma_{t}({\bm{x}})$ with high probability, where $\beta_{t}\to\infty$ as $t\to\infty$. Applying such a $\beta_{t}$ to the asymptotic analysis leads to $c_{\tau}(\beta_{t})$ that increases exponentially with $\sqrt{\beta_{t}}$ as $t\to\infty$. Hence, it is difficult to prove the boundedness of $r_{t}$. The second challenge is the loss of the inequalities $y^{+}_{t}\leq f({\bm{x}}_{t})$ and $f^{+}_{t+1}\leq f^{+}_{t}$. We overcome these challenges by adopting the GP prior assumption, which allows for a pointwise prediction error inequality, i.e., Lemma 3.12 at a given ${\bm{x}}\in C$ and $t$. Thus, $\beta$ is based on the probability $\delta$ and does not increase with $t$. In addition, we use the monotonicity of best observations, i.e., $y^{+}_{t}-y^{+}_{t+1}\geq 0$, while the i.i.d. Gaussian noise gives a probability lower bound where $y^{+}_{i}\leq f({\bm{x}}_{i}),i\in A_{k},|A_{k}|=k$ for a given $k$.
4.2 Improved error bound
An improved error bound compared to Theorem 4.2 is given next by analyzing and utilizing the exploration and exploitation properties of EI under both noiseless and noisy cases (see remark 4.16 for examples of constants).
Theorem 4.7.
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{9c_{\alpha}}{\delta})$, $w=\sqrt{2\log(\frac{9}{2\delta})}$, $c_{t}^{\sigma}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$, and $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Choose the constants $C_{1}=\frac{1}{\Phi(-w)}$ and $C_{2}=\frac{\phi(0)}{\Phi(-w)}+\sqrt{\beta}$. The EI algorithm leads to the error bound
(70) $\mathbb{P}\left\{r_{t}\leq 2C_{1}(M+\sqrt{c_{t}^{\sigma}}\sigma)\frac{3}{t-3}+(C_{1}\sqrt{\beta}+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right\}\geq 1-\delta,$
for $t\geq\frac{3\log(\frac{3}{\delta})}{\log(2)}+3$ and some $t_{k}\in[\frac{t}{3}-1,t]$. If there is no noise, let $\beta=2\log(\frac{3c_{\alpha}}{\delta})$, $C_{1}=\frac{1}{\Phi(-\sqrt{\beta})}$, $C_{2}=\frac{\phi(0)}{\Phi(-\sqrt{\beta})}+\sqrt{\beta}$. Then,
(71) $\mathbb{P}\left\{r_{t}\leq 2C_{1}M\frac{2}{t-2}+(C_{1}\sqrt{\beta}+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right\}\geq 1-\delta,$
where $t_{k}\in[\frac{t}{2}-1,t]$.
Proof 4.8.
First, notice that the choice of $\beta$ satisfies $\beta>1.44$ and thus is larger than the values required in Lemma 3.15 and Lemma 3.12 for $\frac{\delta}{9}$, i.e., (33) and (28) stand with probability greater than $1-\frac{\delta}{9}$ each. Similar to Theorem 4.2, we know that with probability greater than $1-\frac{\delta}{3}$, (63) stands. Therefore, $y^{+}_{t}-y^{+}_{t+1}\geq\frac{2M+2\sqrt{c_{t}^{\sigma}\sigma}}{k}$ at most $k$ times for any $k\in\mathbb{N}$.
Given a sequence of iterates ${\bm{x}}_{i}$ of size $k\in\mathbb{N}$ where $i\in A_{k}$, $A_{k}$ being the index set of samples, the probability of $f({\bm{x}}_{i})<y_{i}=f({\bm{x}}_{i})+\epsilon({\bm{x}}_{i})$ for all $i\in A_{k}$ is $(\frac{1}{2})^{k}$. Let $k=[\frac{t}{3}]$ so that $3k\leq t\leq 3(k+1)$. Consider now $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$, which implies $(\frac{1}{2})^{k}\leq(\frac{1}{2})^{\frac{t}{3}-1}\leq\frac{\delta}{3}$. Thus, there exists $k\leq t_{k}\leq 3k$ so that
(72) $\mathbb{P}\left\{y_{t_{k}}^{+}-y_{t_{k}+1}^{+}<\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k},\ f({\bm{x}}_{t_{k}+1})\geq y_{t_{k}+1}\geq y_{t_{k}+1}^{+}\right\}\geq 1-\frac{2\delta}{3},$
for $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$. Let $t\geq\frac{3}{\log(2)}\log(\frac{3}{\delta})+3$.
From the monotonicity of $y_{t}^{+}$, (8) and Lemma 3.15,
(73) $r_{t}=$ $y^{+}_{t}-f({\bm{x}}^{*})\leq y^{+}_{t_{k}}-f({\bm{x}}^{*})\leq I_{t_{k}}({\bm{x}}^{*})$
$\leq$ $EI_{t_{k}}({\bm{x}}^{*})+\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}^{*})\leq EI_{t_{k}}({\bm{x}}_{t_{k}+1})+\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}^{*})$
$=$ $(y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1}))\Phi(z_{t_{k}}({\bm{x}}_{t_{k}+1}))+\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\phi(z_{t_{k}}({\bm{x}}_{t_{k}+1}))+\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}^{*})$
$\leq$ $(y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1}))\Phi(z_{t_{k}}({\bm{x}}_{t_{k}+1}))+\phi(0)\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})+\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}^{*}),$
with probability greater than $1-\frac{\delta}{9}$, where the last inequality uses Lemma 3.1. From Lemma 3.12,
(74) $\mathbb{P}\left\{|f({\bm{x}}_{t_{k}+1})-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})|\leq\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right\}\geq 1-\frac{\delta}{9}.$
From (72) and (74), with probability greater than $1-\frac{7\delta}{9}$,
(75) $y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})=$ $y^{+}_{t_{k}}-y^{+}_{t_{k}+1}+y^{+}_{t_{k}+1}-f({\bm{x}}_{t_{k}+1})+f({\bm{x}}_{t_{k}+1})-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})$
$<$ $\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k}+\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}).$
Applying (75) to (73), we have with probability greater than $1-\frac{8\delta}{9}$,
(76) $r_{t}\leq$ $\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k}+(\phi(0)+\sqrt{\beta})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})+\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}^{*}).$
From (76), in order to obtain the upper bound for $r_{t}$, we need to consider the convergence behavior of $\sigma_{t_{k}}({\bm{x}}^{*})$ with respect to $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$. By the definition of ${\bm{x}}_{t_{k}+1}$, it generates the largest $EI_{t_{k}}$, i.e.,
(77) $EI_{t_{k}}({\bm{x}}_{t_{k}+1})\geq EI_{t_{k}}({\bm{x}}^{*}).$
We aim to show that (77) guarantees a small $\sigma_{t_{k}}({\bm{x}}^{*})$ in some scenarios.
For simplicity of presentation, define $w=\sqrt{2\log(\frac{9}{2\delta})}$. From the proof of Lemma 3.12 and Lemma 3.1, we know for $w>0$,
(78) $\mathbb{P}\{f({\bm{x}}^{*})-\mu_{t_{k}}({\bm{x}}^{*})>-\sigma_{t_{k}}({\bm{x}}^{*})w\}=1-\Phi(-w)\geq 1-\frac{1}{2}e^{-\frac{w^{2}}{2}}=1-\frac{\delta}{9}.$
Then, with probability greater than $1-\frac{\delta}{9}$, $f({\bm{x}}^{*})-\mu_{t_{k}}({\bm{x}}^{*})>-w\sigma_{t_{k}}({\bm{x}}^{*})$. By (74) and (78), with probability greater than $1-\frac{2\delta}{9}$,
(79) $y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}^{*})>y^{+}_{t_{k}}-f({\bm{x}}^{*})-w\sigma_{t_{k}}({\bm{x}}^{*})$
$=y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})+\mu_{t_{k}}({\bm{x}}_{t_{k}+1})-f({\bm{x}}_{t_{k}+1})+f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})-w\sigma_{t_{k}}({\bm{x}}^{*})$
$\geq y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})+f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})-w\sigma_{t_{k}}({\bm{x}}^{*})-\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}).$
Notice that (76) and (79) come from (72), (73), (74) and (78). Therefore, by union bound, (76) and (79) both stand with probability greater than $1-\delta$. By the definition of $C_{1}$ and $C_{2}$, and Lemma 3.1,
(80) $C_{1}=\frac{1}{\Phi(-w)}>2,\ C_{2}-\sqrt{\beta}=\frac{\phi(0)}{\Phi(-w)}>\frac{\phi(0)}{\frac{1}{2}e^{-\frac{w^{2}}{2}}}=2\phi(0)e^{\frac{w^{2}}{2}}>w+1,$
where the last inequality uses $w\geq\sqrt{2\log(\frac{9}{2})}$. We consider two scenarios regarding the term $f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})$ in (79).
Scenario A. Suppose first that $f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})$ satisfies
(81) $f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})\leq C_{1}\max\{y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1}),0\}+C_{2}\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}).$
From the monotonicity of $y_{t}^{+}$, (75) and (81), we have with probability $\geq 1-\frac{7\delta}{9}$,
(82) $r_{t}=$ $y^{+}_{t}-f({\bm{x}}^{*})\leq y^{+}_{t_{k}+1}-f({\bm{x}}^{*})\leq f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})$
$\leq$ $C_{1}\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k}+(C_{1}\sqrt{\beta}+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}),$
which then proves (70). Notice that $y_{t_{k}+1}^{+}\leq f({\bm{x}}_{t_{k}+1})$ is used again in (82).
Scenario B. Next, we analyze the B scenario where
(83) $f({\bm{x}}_{t_{k}+1})-f({\bm{x}}^{*})>C_{1}\max\{y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1}),0\}+C_{2}\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}).$
Since the exploitation part $y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})$ is bounded by $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$ in (75) and thus decreases for large $t$, our idea is to show that by choosing finite $C_{1}$ and $C_{2}$ large enough, EI would focus on exploration and (77) would ensure $\sigma_{t_{k}}({\bm{x}}^{*})$ decreases. Specifically, $\sigma_{t_{k}}({\bm{x}}^{*})\leq\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$ with a high probability under scenario B (83). An illustration of this idea is shown in Figure 2. We prove by contradiction. Suppose on the contrary,
(84) $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})<\sigma_{t_{k}}({\bm{x}}^{*})\leq 1,$
with probability $1-\delta_{\sigma}$.
To proceed, we use the notation of exploitation and exploration for EI (6). Let $a_{t_{k}}=y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})$ and $b_{t_{k}}=\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$, i.e., $EI(a_{t_{k}},b_{t_{k}})=EI_{t_{k}}({\bm{x}}_{t_{k}+1})$. Based on (79) and (84), let $b^{*}_{t_{k}}=\sigma_{t_{k}}({\bm{x}}^{*})$ and
(85) $a^{*}_{t_{k}}$ $=y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1})+C_{1}\max\{y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}_{t_{k}+1}),0\}+C_{2}\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$
$-w\sigma_{t_{k}}({\bm{x}}^{*})-\sqrt{\beta}\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})=a_{t_{k}}+C_{1}\max\{a_{t_{k}},0\}-wb^{*}_{t_{k}}+(C_{2}-\sqrt{\beta})b_{t_{k}}.$
Thus, $b^{*}_{t_{k}}>b_{t_{k}}$ with probability $1-\delta_{\sigma}$ by (84). Moreover, using (79), (83), and (85)
(86) $y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}^{*})\geq a^{*}_{t_{k}},$
with probability $\geq 1-\frac{2\delta}{9}$. Since EI is monotonically increasing with respect to both $a$ and $b$ by Lemma 3.8, from (86), we have
(87) $EI_{t_{k}}({\bm{x}}^{*})=EI(y^{+}_{t_{k}}-\mu_{t_{k}}({\bm{x}}^{*}),\sigma_{t_{k}}({\bm{x}}^{*}))\geq EI(a^{*}_{t_{k}},b^{*}_{t_{k}}),$
with probability $\geq 1-\frac{2\delta}{9}$. Denote $C_{3}=C_{2}-\sqrt{\beta}$, where by (80) $C_{3}>w+1$. We consider two cases regarding $b_{t_{k}}^{*}$ in (85).
Case 1. If
(88) $b^{*}_{t_{k}}<\frac{C_{1}}{w}\max\{a_{t_{k}},0\}+\frac{C_{3}}{w}b_{t_{k}},$
then by (85), $a^{*}_{t_{k}}>a_{t_{k}}$. Further, from (84), with probability $1-\delta_{\sigma}$,
(89) $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})>EI(a_{t_{k}},b_{t_{k}})=EI_{t_{k}}({\bm{x}}_{t_{k}+1}).$
Case 2. Next we consider the case where
(90) $b^{*}_{t_{k}}\geq\frac{C_{1}}{w}\max\{a_{t_{k}},0\}+\frac{C_{3}}{w}b_{t_{k}}\geq\left(\frac{C_{3}}{w}\right)b_{t_{k}}.$
From $\frac{C_{3}}{w}>1$, it is clear that in Case 2., (84) always holds. We further distinguish between two cases based on the sign of $a_{t_{k}}$.
Case 2.1. Consider $a_{t_{k}}<0$ and $z_{t_{k}}=\frac{a_{t_{k}}}{b_{t_{k}}}<0$. Using (85) and (90),
(91) $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})=$ $EI(a_{t_{k}}+C_{3}b_{t_{k}}-wb^{*}_{t_{k}},b^{*}_{t_{k}})=b^{*}_{t_{k}}\tau\left(\frac{a_{t_{k}}+C_{3}b_{t_{k}}-wb^{*}_{t_{k}}}{b^{*}_{t_{k}}}\right)$
$=$ $\frac{1}{\rho_{t_{k}}}b_{t_{k}}\tau\left((z_{t_{k}}+C_{3})\rho_{t_{k}}-w\right).$
where $\rho_{t_{k}}=\frac{b_{t_{k}}}{b^{*}_{t_{k}}}$. Consider two cases based on the value of $z_{t_{k}}$.
Case 2.1.1. Suppose $z_{t_{k}}>-C_{3}$. Define function $\tilde{\tau}:\mathbb{R}\to\mathbb{R}$ at $z,w$ as
(92) $\bar{\tau}(\rho;z,w)=\frac{1}{\rho}\tau\left((z+C_{3})\rho-w\right),$
where under the conditions of Case 2. and Case 2.1.1., $\rho\in(0,\frac{w}{C_{3}}),-C_{3}<z<0$. We aim to find the minimum of (92) given $z$ and $w$. Taking the derivative with $\rho$ using Lemma 3.2, we have
[…]
Define the function $\theta(\rho;z,w)=-w\Phi\left((z+C_{3})\rho-w\right)+\phi((z+C_{3})\rho-w)$. From (93), the sign of $\frac{d\bar{\tau}}{d\rho}$ is determined by $\theta(\rho;z,w)$. The derivative of $\theta$ with $\rho$ is
(94) $\frac{d\theta}{d\rho}=$ $-w\phi((z+C_{3})\rho-w)(z+C_{3})-\phi((z+C_{3})\rho-w)((z+C_{3})\rho-w)(z+C_{3})$
$=$ $-\phi((z+C_{3})\rho-w)(z+C_{3})^{2}\rho<0,$
for $\forall\rho\in(0,\frac{w}{C_{3}})$. Therefore, $\theta$ is monotonically decreasing with $\rho$. Now let $\rho\to 0$. We have $\theta(\rho;z,w)\to-w\Phi(-w)+\phi(-w)=\tau(-w)>0$. Further, let $\rho\to\frac{w}{C_{3}}$. Then, $\theta(\rho;z,w)\to-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)$. Depending on the sign of $-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)$, $\bar{\tau}$ has two different types of behavior.
Case 2.1.1.1 First, $-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)\geq 0$, which means $\theta(\rho;z,w)>0$ for $\forall\rho\in(0,\frac{w}{C_{3}})$. By (93), $\bar{\tau}(\rho;z,w)$ is monotonically decreasing with $\rho$ and
(95) $\bar{\tau}(\rho;z,w)>\bar{\tau}(\frac{w}{C_{3}};z,w)=\frac{C_{3}}{w}\tau(\frac{w}{C_{3}}z).$
Using $\frac{w}{C_{3}}<1$ and the monotonicity of $\tau$ (Lemma 3.2), we have
(96) $\bar{\tau}(\rho;z,w)>\frac{C_{3}}{w}\tau(\frac{w}{C_{3}}z)>\tau(\frac{w}{C_{3}}z)>\tau(z).$
Applying (96) to (91) with $z=z_{t_{k}}$,
(97) $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})=b_{t_{k}}\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)>b_{t_{k}}\tau(z_{t_{k}})=EI(a_{t_{k}},b_{t_{k}}).$
Case 2.1.1.2 In the second case, $-w\Phi(\frac{w}{C_{3}}z)+\phi(\frac{w}{C_{3}}z)<0$, which means there exists a unique $\bar{\rho}\in(0,\frac{w}{C_{3}})$ as a stationary point for $\theta(\rho;z,w)$, i.e.,
(98) $-w\Phi((z+C_{3})\bar{\rho}-w)+\phi((z+C_{3})\bar{\rho}-w)=0.$
We claim $\bar{\rho}$ is a local minimum of $\bar{\tau}$ via the second derivative test. From (93), (94), and (98), the second-order derivative of $\bar{\tau}$ with $\rho$ at $\bar{\rho}$ is
(99) $\frac{d^{2}\bar{\tau}}{d\rho^{2}}|_{\bar{\rho}}=$ $\frac{2}{\bar{\rho}^{3}}\theta(\bar{\rho};z,w)-\frac{1}{\bar{\rho}^{2}}\frac{d\theta}{d\rho}(\bar{\rho};z,w)=\frac{1}{\rho}\phi((z+C_{3})\bar{\rho}-w)(z+C_{3})^{2}>0.$
It is well-known from second derivative test that $\bar{\rho}$ is a local minimum of $\bar{\tau}$. Further, $\frac{d\bar{\tau}}{d\rho}<0$ for $\rho<\bar{\rho}$ and $\frac{d\bar{\tau}}{d\rho}>0$ for $\rho>\bar{\rho}$. Therefore, $\bar{\rho}$ is the global minimum on $(0,\frac{w}{C_{3}})$ and by (98),
(100) $\bar{\tau}(\bar{\rho};z,w)$ $=\frac{1}{\bar{\rho}}[((z+C_{3})\bar{\rho}-w)\Phi\left((z+C_{3})\bar{\rho}-w\right)+\phi\left((z+C_{3})\bar{\rho}-w\right)]$
$=\frac{1}{\bar{\rho}}[((z+C_{3})\bar{\rho}-w)\Phi\left((z+C_{3})\bar{\rho}-w\right)+w\Phi\left((z+C_{3})\bar{\rho}-w\right)]$
$=(z+C_{3})\Phi\left((z+C_{3})\bar{\rho}-w\right).$
Denote the corresponding $\bar{\rho}$ for $z_{t_{k}}$ as $\bar{\rho}_{t_{k}}$. By (100),
(101) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)-\tau(z_{t_{k}})>$ $(z_{t_{k}}+C_{3})\Phi\left((z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w\right)-\tau(z_{t_{k}})$
If $z_{t_{k}}\in(\frac{C_{3}\bar{\rho}_{t_{k}}-w}{1-\bar{\rho}_{t_{k}}},0)$, we know $(z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w<z_{t_{k}}$. By $\phi(z_{t_{k}})<\phi(0)$, (101) becomes
(102) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)-\tau(z_{t_{k}})>$ $(z_{t_{k}}+C_{3})\Phi\left((z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w\right)-z_{t_{k}}\Phi(z_{t_{k}})-\phi(0)$
$\geq z_{t_{k}}[\Phi((z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}$ $-w)-\Phi(z_{t_{k}})]+C_{3}\Phi\left((z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w\right)-\phi(0)$
$\geq C_{3}\Phi((z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w)$ $-\phi(0)>C_{3}\Phi(-w)-\phi(0)=C_{3}\Phi(-w)-\phi(0)\geq 0,$
where the last inequality is from (80). If $z_{t_{k}}\leq\frac{C_{3}\bar{\rho}_{t_{k}}-w}{1-\bar{\rho}_{t_{k}}}$, we have $(z_{t_{k}}+C_{3})\bar{\rho}_{t_{k}}-w\geq z_{t_{k}}$. Suppose now $z_{t_{k}}>-C_{3}+1$. By (101) and Lemma 3.4,
(103) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)-\tau(z_{t_{k}})\geq$ $(z_{t_{k}}+C_{3})\Phi\left(z_{t_{k}}\right)-\tau(z_{t_{k}})>\Phi(z_{t_{k}})-\tau(z_{t_{k}})>0.$
Finally, if $z_{t_{k}}\leq-C_{3}+1$, from (80), $z_{t_{k}}<-w$. By definition (92),
(104) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)-\tau(z_{t_{k}})\geq$ $\frac{C_{3}}{w}\tau\left(-w\right)-\tau(z_{t_{k}})>\frac{C_{3}}{w}\tau\left(-w\right)-\tau(-w)>0.$
Combining all three cases (102), (103) and (104), (101) leads to
(105) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)-\tau(z_{t_{k}})>0.$
From (91), (92) and (105), we again have (97). Therefore, (97) is valid for Case 2.1.1..
Case 2.1.2. Next, consider $z_{t_{k}}\leq-C_{3}$. By (90), $\frac{b_{t_{k}}}{b^{*}_{t_{k}}}\leq\frac{w}{C_{3}}<1$. Using the monotonicity of $\tau(\cdot)$, (91) implies
(106) $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})=\frac{1}{\rho_{t_{k}}}b_{t_{k}}\tau\left(\rho_{t_{k}}(z_{t_{k}}+C_{3})-w\right)>$ $\frac{C_{3}}{w}b_{t_{k}}\tau\left(\frac{w}{C_{3}}(z_{t_{k}}+C_{3})-w\right)$
$>$ $b_{t_{k}}\tau\left(z_{t_{k}}\right)=EI(a_{t_{k}},b_{t_{k}}).$
That is, we have (97) for Case 2.1.2. Combining Case 2.1.1 and Case 2.1.2, we have $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})>EI(a_{t_{k}},b_{t_{k}})$ if $a_{t_{k}}<0$ (Case 2.1.). Figure 3 demonstrates an example contour of $\bar{\tau}(\rho;z,w)-\tau(z)$ for given $w$ and $C_{3}$.
Case 2.2. Next, we consider $a_{t_{k}}\geq 0$. By definition (85),
(107) $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})=$ $EI((C_{1}+1)a_{t_{k}}+C_{3}b_{t_{k}}-wb^{*}_{t_{k}},b^{*}_{t_{k}})\geq b^{*}_{t_{k}}\tau\left((C_{1}z_{t_{k}}+C_{3})\rho_{t_{k}}-w\right)$
$=$ $b_{t_{k}}\frac{1}{\rho_{t_{k}}}\tau\left(C_{1}z_{t_{k}}\rho_{t_{k}}+C_{3}\rho_{t_{k}}-w\right),$
where $\rho_{t_{k}}=\frac{b_{t_{k}}}{b^{*}_{t_{k}}}$. The inequality in (107) uses $(C_{1}+1)a_{t_{k}}\geq C_{1}a_{t_{k}}$ and the monotonicity of $\tau(\cdot)$. By (90) and $a_{t_{k}}\geq 0$,
(108) $\rho_{t_{k}}\leq\left(\frac{w}{C_{3}}-\frac{C_{1}}{C_{3}}\frac{a_{t_{k}}}{b_{t_{k}}^{*}}\right)\leq\left(\frac{w}{C_{3}}-\frac{C_{1}}{C_{3}}a_{t_{k}}\right)\leq\frac{w}{C_{3}}.$
Thus, $C_{3}\rho_{t_{k}}-w\leq 0$. Define function $\tilde{\tau}:\mathbb{R}^{2}\to\mathbb{R}$ as
(109) $\tilde{\tau}(\rho,z;w)=\frac{1}{\rho}\tau\left(C_{1}z\rho+C_{3}\rho-w\right),$
for $\rho\in(0,\frac{w}{C_{3}}),z\geq 0$. From (107), $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})\geq b_{t_{k}}\tilde{\tau}(\rho_{t_{k}},z_{t_{k}};w)$. We now compare $\tilde{\tau}(\rho_{t_{k}},z_{t_{k}};w)$ and $\tau(z_{t_{k}})$. For $\forall\rho_{t_{k}}\in(0,\frac{w}{C_{3}})$, (80) and Lemma 3.2 lead to the derivative
(110) $\frac{\partial\tilde{\tau}}{\partial z}=C_{1}\Phi\left(C_{1}z\rho+C_{3}\rho-w\right)>C_{1}\Phi(-w)>\Phi(z)=\frac{d\tau}{dz}.$
Therefore, both $\tilde{\tau}(\cdot,\cdot)$ and $\tau(\cdot)$ are monotonically increasing with $z>0$ and $\tilde{\tau}$ always has a larger positive derivative. For any $\rho_{t_{k}}$, this leads to
(111) $\tilde{\tau}(\rho_{t_{k}},z_{t_{k}};w)-\tau(z_{t_{k}})\geq\tilde{\tau}(\rho_{t_{k}},0;w)-\tau(0)=\frac{1}{\rho_{t_{k}}}\tau\left(C_{3}\rho_{t_{k}}-w\right)-\tau(0).$
Notice that the right-hand side of (111) can be written via (92) as
(112) $\bar{\tau}(\rho_{t_{k}};0,w)-\tau(0).$
Therefore, we can follow the analysis in Case 2.1.1 with $z=0$. Notice that $-w\Phi(0)+\phi(0)<0$ for $\forall w>1$. Thus, (112) follows Case 2.1.1.2 where a minimum exists at $\bar{\rho}\in(0,\frac{w}{C_{3}})$ which satisfies $w\Phi(C_{3}\bar{\rho}-w)=\phi(C_{3}\bar{\rho}-w)$. Notice that $\bar{\rho}$ does not depend on $t_{k}$ anymore. Hence, we have
[…]
Given (80),
(114) $\bar{\tau}(\rho_{t_{k}};0,w)-\tau(0)>C_{3}\Phi(-w)-\phi(0)>0.$
By (107), (109), (111), (112), and (114), we have
(115) $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})-EI(a_{t_{k}},b_{t_{k}})\geq b_{t_{k}}\tilde{\tau}(\rho_{t_{k}},0;w)-b_{t_{k}}\tau(0)>0.$
Illustrative examples of $\tilde{\tau}$ is given in Figure 4.
To summarize, under the Scenario B Case 2, for all cases given (90),
$EI(a_{t_{k}}^{*},b_{t_{k}}^{*})>EI(a_{t_{k}},b_{t_{k}})$. Combined with (89), under Scenario B, $EI(a_{t_{k}}^{*},b_{t_{k}}^{*})>EI(a_{t_{k}},b_{t_{k}})$ with probability greater than $1-\delta_{\sigma}$. Combined with (87), with probability greater than $1-\frac{2\delta}{9}-\delta_{\sigma}$, $EI_{t_{k}}({\bm{x}}^{*})>EI_{t_{k}}({\bm{x}}_{t_{k}+1})$. However, this is a contradiction of (77), a sure event. Therefore, we have
(116) $1-\frac{2\delta}{9}-\delta_{\sigma}\leq 0$
That is, $\delta_{\sigma}\geq 1-\frac{2\delta}{9}$. From (84),
(117) $\mathbb{P}\{\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\geq\sigma_{t_{k}}({\bm{x}}^{*})\}\geq 1-\frac{2\delta}{9}.$
Notice again that (117) and (76) both stand with probability greater than $1-\delta$, as $\delta$ is derived from (72), Lemma 3.15 and Lemma 3.12. Under Scenario B, by (117) and (76), with probability greater than $1-\delta$,
(118) $r_{t}\leq\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k}+(\phi(0)+2\sqrt{\beta})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}).$
From (80), $C_{1}>2$ and $C_{2}>1$. Combining Scenario A (82) and Scenario B (118), we obtain (70).
If there is no noise, i.e., $\sigma=0$, we have $f({\bm{x}}_{t})=y_{t}\geq y_{t}^{+}$. Therefore, we can choose $k=[\frac{t}{2}]$ and there exists $k\leq t_{k}\leq 2k$ so that $y_{t_{k}}^{+}-y_{t_{k}+1}^{+}<\frac{2M}{k}$ with probability $1$. Following the same proof procedure as the noisy scenario, we obtain (71).
4.3 Convergence rate and comparative analysis
The rate of decrease for the noiseless error bound of Theorem 4.7 is given below in the setup of [14]. The readers are referred to [14] for the underlying assumptions on the kernels and the choice of parameters $\nu>0$ and $\alpha\geq 0$.
Theorem 4.9.
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{3c_{\alpha}}{\delta})$, where $c_{\alpha}=\frac{1+2\pi}{\pi}$, $C_{1}=\frac{1}{\Phi(-\sqrt{\beta})}$ and $C_{2}=\frac{\phi(0)}{\Phi(-\sqrt{\beta})}+\sqrt{\beta}$. Suppose the kernels $k(\cdot,\cdot)$ satisfy assumptions 1-4 in [14], then there exists constant $C^{\prime}$ so that
(119) $\mathbb{P}\left\{r_{t}\leq 2C_{1}M\frac{3}{t-3}+(C_{1}\sqrt{\beta}+C_{2})C^{\prime}(\frac{t-3}{3})^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}\left(\frac{t}{3}\right)\right\}\geq 1-\delta.$
Hence, the convergence rate of GP-EI is $\mathcal{O}(t^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(t))$.
Proof 4.10.
By Lemma 7 in [14], there exists a constant $C^{\prime}>0$ so that for $\forall k\in\mathbb{N}$ and ${\bm{x}}_{i},i=1,\dots,t$, $\sigma_{i}({\bm{x}}_{i+1})\geq C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)$ holds at most $k$ times, where $\eta=\alpha$ if $\nu\leq 1$ and $\eta=0$ if $\nu>1$. Both $\nu>0$ and $\alpha\geq 0$ are parameters defining the properties of the kernels in assumptions 1-4 in [14].
Following the proof of Theorem 4.7, we can let $k=\left[\frac{t}{3}\right]$. Then, there exists $k\leq t_{k}\leq 3k$ so that $y_{t_{k}}^{+}-y_{t_{k}+1}<\frac{2M}{k}$ and $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})<C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)$. Applying Theorem 4.7, with probability $1-\delta$,
(120) $r_{t}\leq$ $2C_{1}M\frac{1}{k}+(C_{1}\sqrt{\beta}+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$
$\leq$ $2C_{1}M\frac{1}{k}+(C_{1}\sqrt{\beta}+C_{2})C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k).$
Remark 4.11.
Theorem 4.9 illustrates that the convergence rate in the noiseless case under the GP prior assumption is similar to that under the RKHS assumption in [14]. The analysis can be applied to Theorem 4.2 in the noiseless case as well, with a similar convergence rate.
The rate of decrease for the error bound with noise is given below.
Theorem 4.12.
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{9c_{\alpha}}{\delta})$, $w=\sqrt{2\log(\frac{9}{2\delta})}$, and $c_{t}^{\sigma}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$, where $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Let $C_{1}=\frac{1}{\Phi(-w)}$ and $C_{2}=\frac{\phi(0)}{\Phi(-w)}+\sqrt{\beta}$. Then, for $t\geq\frac{4\log(\frac{3}{\delta})}{\log(2)}+4$, the error bound of the EI algorithm reduces at the rate
(121) $\mathcal{O}\left(t^{-\frac{1}{2}}\log(t)^{\frac{d+1}{2}}\right)\ \text{and }\ \mathcal{O}(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)),$
with probability greater than $1-\delta$, for SE and Matérn kernels, respectively.
Proof 4.13.
From Lemma 2.1, we know $\sum_{i=0}^{t-1}\sigma_{i}^{2}({\bm{x}}_{i+1})\leq C_{\gamma}\gamma_{t}$, where $C_{\gamma}=\frac{2}{\log(1+\sigma^{-2})}$. Therefore, $\sigma_{i}^{2}({\bm{x}}_{i+1})\geq\frac{C_{\gamma}\gamma_{t}}{k}$ at most $k$ times for any $k\in\mathbb{N}$. Choose $k=[\frac{t}{4}]$ so that $4k\leq t\leq 4(k+1)$. Then, for $t\geq\frac{4\log(\frac{3}{\delta})}{\log(2)}+4$, $(\frac{1}{2})^{k}\leq(\frac{1}{2})^{\frac{t}{4}-1}\leq\frac{\delta}{3}$. Following the proof of Theorem 4.7, there exists $k\leq t_{k}\leq 4k$ where $y^{+}_{t_{k}}-y^{+}_{t_{k}+1}<\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{k}$, $f({\bm{x}}_{t_{k}+1})\geq y_{t_{k}+1}$, and $\sigma_{t_{k}}^{2}({\bm{x}}_{t_{k}+1})<\frac{C_{\gamma}\gamma_{t}}{k}$, with probability greater than $1-\frac{2\delta}{3}$ for $t\geq\frac{4\log(\frac{3}{\delta})}{\log(2)}+4$. Using Theorem 4.7, we can obtain
(122) $r_{t}\leq$ $C_{1}(2M+2\sqrt{c_{t}^{\sigma}}\sigma)\frac{4}{t-4}+(C_{1}\sqrt{\beta}+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})$
$\leq$ $C_{1}(2M+2\sqrt{c_{t}^{\sigma}}\sigma)\frac{4}{t-4}+2(C_{1}\sqrt{\beta}+C_{2})\sqrt{\frac{C_{\gamma}\gamma_{t}}{t-4}},$
with probability greater than $1-\delta$. If $k(\cdot,\cdot)$ is the SE kernel, from Lemma 2.2, the upper bound for $r_{t}$ is $\mathcal{O}(t^{-\frac{1}{2}}(\log t)^{\frac{d+1}{2}})$. Similarly, by Lemma 2.2 and (122), the rate of reduction of $r_{t}$ for Matérn kernel can be obtained as well.
Next, we prove the error bound in Theorem 4.7 is smaller than that of Theorem 4.2.
Proposition 4.14.
Using the same $\delta\in(0,1)$ in Theorem 4.2 and 4.7, the error bound in Theorem 4.7 is smaller than that of Theorem 4.2.
Proof 4.15.
We prove the noisy case as the noiseless case is similar. The corresponding $\beta_{2}$ in Theorem 4.7 is $\beta_{2}=2\log(\frac{9c_{\alpha}}{\delta})$, and the $\beta_{1}$ in Theorem 4.2 is $\beta_{1}=2\log(\frac{6}{\delta})$. Clearly, the probabilities for (60) and (70) are the same. To better compare the two theorems, we can rewrite both  (60) and (70) as
(123) $r_{t}\leq 3C_{4}\frac{2M+2\sqrt{c_{t}^{\sigma}}\sigma}{t-3}+C_{5}\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}).$
The bounds in (60) leads to
(124) $C_{4}^{1}=\frac{\tau(\sqrt{\beta_{1}})}{\tau(-\sqrt{\beta_{1}})},\ C_{5}^{1}=\frac{\tau(\sqrt{\beta_{1}})}{\tau(-\sqrt{\beta_{1}})}(\sqrt{\beta_{1}}+\phi(0)),$
where the superscript $1$ indicates Theorem 4.2 and $2$ refers to Theorem 4.7. Let $w_{2}=2\log(\frac{9}{2\delta})$. From (70),
(125) $C_{4}^{2}=\frac{1}{\Phi(-\sqrt{w_{2}})},\ C_{5}^{2}=\frac{\sqrt{\beta_{2}}}{\Phi(-\sqrt{w_{2}})}+\frac{\phi(0)}{\Phi(-\sqrt{w_{2}})}+\sqrt{\beta_{2}}.$
Using $\beta_{1}>w_{2}$, we have
(126) $C_{4}^{1}=\frac{\tau(\sqrt{\beta_{1}})}{\tau(-\sqrt{\beta_{1}})}>\frac{\tau(\sqrt{\beta_{1}})}{\Phi(-\sqrt{\beta_{1}})}>\frac{\sqrt{\beta_{1}}}{\Phi(-\sqrt{w_{2}})}>C_{4}^{2},$
where the first inequality uses Lemma 3.4, the second 3.6, and the last $\beta_{1}>1$. Notice that $w_{2}>2\log(\frac{9}{2})>3$. Then, using $\delta\in(0,1)$, $w_{2}<\beta_{2}$, and inequalities from (126), we have
(127) $C_{5}^{2}=$ $\frac{1}{\Phi(-\sqrt{w_{2}})}(\sqrt{\beta_{2}}+\phi(0)+\sqrt{\beta_{2}}\Phi(-\sqrt{w_{2}}))<\frac{1}{\Phi(-\sqrt{\beta_{1}})}(1.042\sqrt{\beta_{2}}+\phi(0)).$
It is easy to verify that $1.042\sqrt{\beta_{2}}+\phi(0)<\beta_{1}+\phi(0)\sqrt{\beta_{1}}$ by taking the square on both sides. Hence, by Lemma 3.4 and 3.6
(128) $C_{5}^{2}<$ $\frac{1}{\Phi(-\sqrt{\beta_{1}})}(\beta_{1}+\phi(0)\sqrt{\beta_{1}})<\frac{\tau(\sqrt{\beta_{1}})(\sqrt{\beta_{1}}+\phi(0))}{\tau(-\sqrt{\beta_{1}})}=C_{5}^{1}.$
Remark 4.16.
An example of Proposition 4.14 is $\delta_{1}=0.1$. Then, $\beta_{1}=8.19$, $C_{4}^{1}=4632$ and $C_{5}^{1}=15103$. Meanwhile, $\beta_{2}=9.17$, $C_{1}=345$, and $C_{2}=141$. Thus, $C_{4}^{2}=345$, and $C_{5}^{2}=1187$. An illustration of Proposition 4.14 is shown in Figure 5.
Remark 4.17.
Another improvement of an upper bound in the form of Theorem 4.7 compared to Theorem 4.2 is that the cumulative regret can be computed with a sum of $\sigma_{t}({\bm{x}}^{*})$. Then, following the proof of Theorem 4.7, the cumulative regret bounds may be established for GP-EI for the first time. We leave this topic for future investigation.
4.4 Improved error bound for RKHS objective
In this section, we demonstrate how the analysis and techniques from Section 4.1 and 4.3 can be applied to RKHS objective to improve existing best-known error bounds and convergence rates from [14] in the noiseless case. Formally, Assumptions 2.52.6 are valid in this section. Additionally, we maintain a small value of $\sigma$, e.g., $10^{-8}$ in computing the posterior mean and variance in (2), even though there is no observation noise. The following lemma can be found in [29].
Lemma 4.18.
In the noiseless scenario, At ${\bm{x}}\in C$ and $t\in\mathbb{N}$,
(129) $|f({\bm{x}})-\mu_{t}({\bm{x}})|\leq B\sigma_{t}({\bm{x}}).$
Next, we state the relationship between $I_{t}$ and $EI_{t}({\bm{x}})$, similar to that of Lemma 3.15.
Lemma 4.19.
At ${\bm{x}}\in C,t\in\mathbb{N}$,
(130) $|I_{t}({\bm{x}})-EI_{t}({\bm{x}})|\leq(B+1)\sigma_{t}({\bm{x}}).$
The proof of Lemma 4.19 can be seen in Lemma 8 in [14] We state the asymptotic convergence theorem under RKHS assumptions from [14] as a lemma.
Lemma 4.20.
The error of noiseless GP-EI satisfies
(131) $r_{t}\leq c_{\tau}(B)\left[2\frac{2M}{t-2}+(B+\phi(0))\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\right],$
where $c_{\tau}(B)=\frac{\tau(B)}{\tau(-B)}$ and $t_{k}\in[\frac{t}{2}-1,t]$.
The proof of Lemma 4.20 is the same as that of Theorem 4.2 with corresponding versions of Lemma 3.17 and 3.12 in RKHS. It is also Theorem 2 from [14]. The improved error bound is given next.
Theorem 4.21.
Let $C_{1}=\frac{1}{\Phi(-B)}$, $C_{2}=B+\frac{\phi(0)}{\Phi(-B)}$. The GP-EI generates error bound
(132) $r_{t}\leq 2C_{1}\frac{2M}{t-2}+(C_{1}B+C_{2})\sigma_{t_{k}}({\bm{x}}_{t_{k}+1}),$
where $t_{k}\in[\frac{t}{2}-1,t]$. Compared to (131), the error bound is smaller.
Proof 4.22.
The proof of (132) is similar to that of Theorem 4.7 by using Lemma 4.18 and 4.19 and replacing $\sqrt{\beta}$ with $B$. Further, the probability $\delta=0$. Details of the proof to arrive at (132) are omitted.
We now show that the error bound in (132) is better than that in (131). Similar to Proposition 4.14, the constants for (131) are
(133) $C_{4}^{1}=\frac{\tau(B)}{\tau(-B)},\ C_{5}^{1}=\frac{\tau(B)}{\tau(-B)}(B+\phi(0)),$
while the constants for (132) are
(134) $C_{4}^{2}=\frac{1}{\Phi(-B)},\ C_{5}^{2}=B+\frac{B+\phi(0)}{\Phi(-B)}.$
Now define
(135) $c_{r}(B)=\frac{\tau(B)(B+\phi(0))}{\Phi(-B)B+B+\phi(0)}<\frac{\tau(B)\Phi(-B)(B+\phi(0))}{\tau(-B)(\Phi(-B)B+B+\phi(0))}=\frac{C_{5}^{1}}{C_{5}^{2}},$
where the inequality uses Lemma 3.4. Similarly, $\frac{C_{4}^{1}}{C_{4}^{2}}>c_{r}(B)$. Therefore,
(136) $r_{t}^{2}<\frac{1}{c_{r}(B)}r_{t}^{1}.$
Notice that $\frac{1}{c_{r}(B)}\to 0$ as $B\to\infty$. That is, significant improvement in error bound is achieved particularly for a large $B$.
Remark 4.23.
Compared to [14], under the same assumption, Theorem 4.21 provides a provably smaller bound (see Proposition 4.14). The convergence rate can be proven if assumptions (1)-(4) for $k({\bm{x}},{\bm{x}}^{\prime})$ in [14] are satisfied, similar to Theorem 4.9.
5 Numerical illustrations
Since the effectiveness and efficiency of GP-EI are well-documented in literature, in this section, we provide numerical illustrations that supplement the theoretical results. First, we show the illustration of $\tau(z)$ and $\Phi(z)$ for Lemma 3.1 and 3.4 in Figure 1.
Figure 1: Left: the relationship between $\Phi(c)$ and $\frac{1}{2}e^{-\frac{1}{2}c^{2}}$ for $c<0$. Right: $\Phi(z)$ v.s. $\tau(z)$ for $z<0$.
Next, we plot the contour of EI based on its exploration and exploitation values (6) in Figure 2.
Figure 2: Contour plot for EI using its exploration and exploitation form (6). Zoomed in view of a small exploitation ($a$) is given on the right. It is clear that EI contains intrinsic trade-off between exploration and exploitation.
Illustrative examples for functions $\bar{\tau}$ and $\tilde{\tau}$ are given next in Figure 3 and Figure 4, respetively. The plots match our analysis in Theorem 4.7.
Figure 3: Left: contour plot for $\log_{10}(\bar{\tau})$ with varying $\rho\in(0,\frac{w}{C_{3}})$ and $z<0$. Here, $w=2$, $C_{1}=44$ and $C_{3}=18$. Right: log-scale comparison of $\bar{\tau}$ with $\tau$ with fixed $z=10^{-3}$. It is clear that $\bar{\tau}(\rho;10^{-3},2)-\tau(10^{-3})>0$ and takes a minimum around $\rho=0.02$.
Figure 4: Left: contour plot for $\log_{10}(\tilde{\tau})$ with varying $\rho\in(0,\frac{w}{C_{3}})$ and $z>0$, which confirms the monotonicity $\tilde{\tau}$ with $z$. Here, $w=3$, $C_{1}=741,$ and $C_{3}=296$. Right: log-scale comparison of $\tilde{\tau}$ with $\tau$ with fixed $z=0$. It is clear that $\tilde{\tau}(\rho;0,2)-\tau(0)>0$.
Finally, we compare the constant parameters in Proposition 4.14 in Figure 5. Given the same probability, it is obvious that the error bound, reflected as the constant parameters, is improved for at least an order of magnitude for most $\delta$ plotted.
Figure 5: The constant parameters $C_{4}^{1}$, $C_{4}^{2}$, $C_{5}^{1}$, and $C_{5}^{2}$ are plotted in log-scale with respect to $\delta$. It is clear that Theorem 4.7 offers an improved bound, often to at least an order of magnitude.
6 Conclusions
This paper contributes to the convergence theories of one of the most widely used Bayesian optimization algorithm, the expected improvement. We extend the asymptotic convergence results of [14] to the Bayesian setting where $f$ is sampled from a GP prior. Further, we establish the first asymptotic convergence results for GP-EI in the noisy case under the GP prior assumption for $f$. Last but not least, we prove an improved error bound by taking advantage of the exploration and exploitation properties of EI.
Acknowledgments
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract 779 DE–AC52–07NA27344. Release number LLNL-JRNL-870576.